### v1

# spacing=nonfrench

Abstract
Generative Artificial Intelligence (GenAI) browser assistants integrate powerful capabilities of GenAI in web browsers to provide rich experiences such as question answering, content summarization, and agentic navigation. These assistants, available today as browser extensions, can not only track detailed browsing activity such as search and click data, but can also autonomously perform tasks such as filling forms, raising significant privacy concerns. It is crucial to understand the design and operation of GenAI browser extensions, including how they collect, store, process, and share user data. To this end, we study their ability to profile users and personalize their responses based on explicit or inferred demographic attributes and interests of users. We perform network traffic analysis and use a novel prompting framework to audit tracking, profiling, and personalization by the ten most popular GenAI browser assistant extensions.
We find that instead of relying on local in-browser models, GenAI browser assistants largely depend on server-side APIs, which can be invoked automatically without explicit user interaction. When invoked, these GenAI browser assistants collect and share webpage content, often the full HTML DOM and sometimes even the user’s form inputs, with their first-party servers. Some GenAI browser assistants also share identifiers and user prompts with third-party trackers such as Google Analytics. The collection and sharing continues even if a webpage contains sensitive information such as health or personal information such as name or social security number entered in a web form. We find that several GenAI browser assistants infer demographic attributes such as age, gender, income, and interests and use this profile—which carries across browsing contexts—to personalize responses. In summary, our work shows that GenAI browser assistants can and do collect personal and sensitive information for profiling and personalization with little to no safeguards.
1 Introduction
With rapid advancements in Natural Language Processing
over the past few years, Large Language Models (LLMs) have grown in scale to incorporate billions of parameters to significantly enhance their contextual understanding and language generation capabilities [29]. As the foundational technology behind many Generative Artificial Intelligence (GenAI) systems, LLMs have played a pivotal role in driving the growth of GenAI across diverse applications such as real-time user assistance, automated content generation, and interactive chatbots. Besides providing a more conversational and personalized search experience due to their search capabilities, they also support agentic capabilities. For instance, they can make a reservation on behalf of the user by performing appropriate searches and filling out the form.
Popular search engines are already integrating LLMs to enhance their search performance – for example, Google Search displays a Gemini-generated overview [42] in response to the user’s search query, while Microsoft Bing’s Copilot uses OpenAI’s model [28]. However, search engines are limited in scope to analyzing a user’s activity on their own platform, restricting their access into user’s broader browsing behavior across the web. To address this limitation, a class of assistants, which we refer to as “GenAI browser assistants” have emerged to leverage the capabilities of generative models in improving a user’s overall browsing on the Internet.
GenAI-based browser assistants are packaged as extensions that can be installed onto a user’s browser, providing them access to everything that a user does in their browser. More specifically, existence in context of an extension improves the ability of these assistants to track a user’s click, websites browsed, content accessed, queries asked, and personal information entered online by storing identifiers to associate user state across websites. The user’s online browsing activities act as a context to browser assistants to build a profile on them. As a result, although the assistants provide a highly personalized browsing experience, this affects user privacy. An online user may inadvertently end up sharing sensitive information regarding themselves, without realizing the resultant consequences. The shared information could be stored on remote servers of the browser assistants or shared with third-parties to target ads to the user.
Despite operational costs reaching as high as \$700,000 per day [22], most LLM platforms do not yet rely on ads for monetization. However, the industry is seeing a shift towards in-house advertising, as observed with Perplexity AI’s sponsored follow-up questions [46]. This raises user privacy concerns similar to those in the existing online advertising ecosystem. With these developments, our web is moving towards an era of in-browser Generative AI – a browser integrated with a local LLM model so as to maintain privacy by retaining the user’s data within the browser and yet guaranteeing the same performance quality as that of a cloud-based LLM model. Chrome’s Built-in AI is in its origin trial and would allow websites and web applications to perform AI-powered tasks within the user’s browser through APIs such as Prompt API [12], eliminating their need to deploy or manage its own AI models. GenAI-based browser assistants aim to provide a similar experience by either directly or indirectly relying on cloud-based LLM models to personalize user’s browsing experience. As a result, it is important to understand the privacy risks that arise from these GenAI-based browser assistants at the cost of user profiling and personalization. Millions of users already use these assistants to personalize their browsing experience on a daily basis. Despite such a large user base, it is unknown how exactly these browser assistants are designed, if, when, what, and how they handle user-dependent information to personalize their responses and the risks they pose to user privacy. To this end, we propose a novel framework to systematically audit GenAI-based browser assistants by addressing the following research questions:
RQ1. How is the architecture of GenAI browser assistants designed? To this end, we qualitatively analyze network traffic activity across different extensions to reason the following design choices: (1) backend model capabilities, (2) response architecture, (3) context restrictions, and (4) response variability.
RQ2. Do GenAI browser assistants collect and share user information in response to their queries? In order to assess privacy risks, it is necessary to analyze how user data is handled. Users may either inadvertently expose personal information to the browser assistants while they browse private spaces on the web; or browser assistants may actively collect user’s browsing related data while they use them in their day-to-day lives. We focus on analyzing network traffic to study both sides of user tracking – (1) implicit and (2) explicit collection and sharing of user data with first-party or third-party entities.
RQ3. Are GenAI browser assistants capable of profiling a user based on their browsing behaviour to personalize its responses? We define profiling as the process of storing and analyzing the collected user data, including their preferences, behaviors, and characteristics, to build a holistic view about the user. While leveraging the profiled information to create an individualized experience is regarded as personalization. We propose a novel prompting framework that – first, leaks and tests for five user attributes – location, age, gender, wealth, and interests for profiling, and second, understands if GenAI browser assistants use leaked information to personalize their responses.
Our research provides the following key contributions with respect to each of the research questions described previously:
  • Architecture. We analyze 10 of the most popular GenAI browser assistants and find that all-but-one uses server-side response generation, while only one operates client-side. We also observed some assistants to automatically invoke response generation based on the user’s search query. Furthermore, 8 assistants were observed to isolate contexts across page navigations, while the remaining two shared user prompt history across different navigation-based browsing contexts.
  • User Tracking. Usage of assistants while browsing through private online spaces resulted in collection of full DOM to partial webpage by different assistants. Merlin was observed to extract even webform inputs, resulting in collection and sharing of medical health records from university health portal (hem.ucdavis.edu), student academic records (from canvas.edu), and SSN (entered on sa.www4.irs.gov/wmr/). Sider and Merlin shared chat and user identifiers with google-analytics.com while TinaMind shared them with analytics.google.com. Merlin also shared user’s raw query with Google analytics, suggesting potential for tracking and retargeting across Google platforms. User’s chat history from past conversations was also shared with first-party servers of four assistants. Moreover, Harpa and Copilot stored the full history in the background service worker’s IndexedDB storage, suggesting its persistence across browsing sessions.
  • Personalization. For search experiments, four extensions (ChatGPT for Google, Copilot, Monica, and Sider) demonstrated profiling for all five user attributes regarding location, age, gender, wealth and interests as well as in-context and out-of-context profiling and personalization prompts. Two of them (Perplexity and TinaMind) didn’t show strong signals of profiling or personalization while Harpa only showed in-context personalization.
2 Background
In this section, we provide background on GenAI systems. In Section 2.1, we describe evolution, functions and challenges of these systems. Next, in Section 2.2, we discuss how GenAI browser assistants work and functionalities they offer.
2.1 Generative AI Systems
The evolution of GenAI. Artificial Intelligence has a history spanning over 50 years, dating back to the 1960s when rule-based systems were first introduced [15]. Over the years, rule-based models were improved using statistical machine learning approaches and multi-layered neural networks that allowed for development of natural language processing applications. Traditional language models relied on sequential processing that made it difficult to capture long-range dependencies in data. The advent of transformer architectures in 2017 [47] addressed this limitation by allowing all elements of the sequence to be processed simultaneously, enabling the model to capture relationships between words regardless of their position in the sequence. This innovation, coupled with massive datasets, improved computational power, and model architecture advancements led to the development of large language models (LLMs) like the generative pre-trained transformer (GPT) [40].
How do GenAI systems work? Generative AI systems, like large language models, function on iteratively predicting the “next” token in a sequence, based on the context provided by previous tokens, to inform its predictions [54]. The underlying transformer-based architecture allows self-attention mechanism to weigh the importance of different words in a sequence differently [47]. When a user performs a query, the model encodes the input into an embedding space and generates a probability distribution over potential next tokens. It selects a token based on this distribution and repeats this process until the output is complete. Following this approach, GenAI systems are able to generate coherent and context-aware text.
Challenges with GenAI systems. Despite their phenomenal generative capabilities, GenAI systems face several challenges that impact their performance, usability, and broader applicability. It is important to take these into account to truly understand how GenAI systems function. The first fundamental issue lies in its “probabilistic” nature. As aforementioned, the prediction of the next token relies on a probabilistic distribution, resulting in different outputs even for the same input, making these genAI systems less predictable than the deterministic ones. This variability is influenced by the “temperature” parameter that controls the randomness of generated output. The value of this parameter typically ranges between 0 and 1 – where higher values result in selection of less probable tokens, either producing more creative responses or outputs that are less coherent or relevant [38]. A lower temperature, on the other hand, results in more deterministic responses. The second challenge is the high cost of training these models. OpenAI’s GPT-4 cost up to \$78M, while Google’s Gemini cost upto \$191M – 15 times higher than the training cost of its precursor models [7]. This limits the frequency at which these models are trained. Moreover, model training leverages large-scale training data to improve its ability to generate more accurate and enhanced responses to user queries. However, it fails to incorporate recent information (e.g., recent news) that was not part of the training process without access to live search capabilities. Lastly, while general-purpose models excel at a wide-range of functions, adapting GenAI systems to domain-specific applications like personalized assistants demands significant effort as well as resources. These standalone models have access to just the user-performed queries on LLM platform’s website limiting their context to only what is explicitly provided by the user in their prompts when they visit their platform website. This makes personalization challenging.
2.2 Generative AI Browser Assistants
How do GenAI browser assistants function? To fully leverage capabilities of GenAI models to personalize user experience on the web, a category of GenAI systems – Gen AI browser assistants have recently emerged. These are essentially browser extensions that function as wrappers on the top of open-source models like OpenAI’s ChatGPT [36], Google’s Gemini [16], Meta’s Llama [26], etc. They record activities that a user performs on the web and queries the open-source model in the backend as per the user’s need to provide a more personalized response. Distinct from the foreground user activities, these browser assistants operate in a separate context as a service worker that is always running in the background when the extension is active. However, extension’s background code can inject Javascript snippets called ‘content scripts’ in every page that a user visits in the foreground. These content scripts can further request and include additional scripts to support necessary extension functionalities such as JQuery for simplified DOM manipulation for instance. Content scripts can also contact background script to share logged user activities. These capabilities provide GenAI browser assistants with kind of a super access to everything that a user does in their browser. For example, assistants can monitor all searches made by the user in their browser, access what content they seek daily on the web, etc. Monitoring user journey across the web, helps browser assistants suggest what a user might be looking for ‘in-the-moment’ by incorporating a highly specific context that is needed for better personalization. For instance, writing assistants can adapt to an individual’s writing style to auto-generate email replies. In summary, GenAI browser assistants address numerous challenges described above related to generic GenAI systems to truly personalize user’s online experience and enhance user interaction and productivity.
Functionalities offered by GenAI Browser Assistants. The need for GenAI browser assistants stems from the growing complexity of digital ecosystems, where users are often overloaded with too much information and require personalized and efficient solutions to manage their needs as they navigate through the web. Based on the offered functionalities, we classify GenAI browser assistants into 3 broad categories – search-based assistants, content-based assistants, and automation assistants. Search-based assistants primarily improve user’s browsing by offering features such as natural language responses to search queries. They also support web page summarization, interactive Q&A with webpages, and text highlighting. Content-based assistants manage content-specific tasks such as writing tasks, social media post generation, meeting transcript summarization, YouTube video explanation, SEO generation, etc. Automation assistants focus on automating workflows to streamline tasks. They include tools that aid in auto-filling forms, coding and debugging agents, task scheduling, automated data extraction and scraping, voice-enabled assistance, etc. All these assistants thus act as intelligent intermediaries to provide a wide array of functionalities to enhance user experience through context-aware suggestions, making them more intuitive, efficient, and tailored to individual needs.
3 Related Works
In this section, we provide literature overview of related works across two dimensions – privacy issues in extensions and privacy issues in GenAI models. We conclude by discussing how our work compares to the prior research in this space.
Privacy Issues in Extensions. Despite enhanced user experience, browser extensions pose significant privacy risks due to their access to sensitive user data, such as browsing history, local storage, payment information, and more [41]. These risks stem from the potential exploitation of extensions by malicious actors seeking unauthorized access to private information [13, 33]. Researchers have examined various aspects of these privacy risks such as detecting spying extensions [3], identifying inconsistencies in privacy practices [8], and finding privacy leaks [50]. There has also been some work to approach privacy issues from security perspective that include detecting vulnerable and malicious extensions [43], cloud-based security analysis [10], and experimental security analysis of sensitive data access [31, 5]. Chen et al. [9] revealed that 2.13% of Chrome browser extensions, representing over 60M users, can leak sensitive information such as browsing history, open tabs, passwords, and location data. Xie et al. [51] further investigated this issue, identifying hundreds of extensions that automatically extract user content from web pages [27]. Notable examples include popular extensions such as Paypal’s Honey, Capital One Shopping, Hola VPN, and Avira Safe Shopping, which exfiltrate browsing URLs and other sensitive information. Alarmingly, approximately 40% of Chrome extensions, including widely used ones, exhibit security vulnerabilities that could compromise private user data [33]. Moreover, Ling et al. [25] discovered that 92% of browser extensions engage in data collection practices that contradict their own privacy policies or stated practices. Many extensions were found to collect excessive data beyond what is necessary for their intended functionality. For instance, while the Chrome extension ‘InserLearning’ claimed to only collect a user’s name, Google account (email address), and Google profile image, it also gathered website content. Additionally, Carnus et al. [20] demonstrated that some extensions leak user data, such as email addresses, names, and phone numbers, to their developers and third-party entities.
Privacy Issues in GenAI models. Past work in this space has mostly focused on discovering privacy issues resulting from the training phase of GenAI models. Researchers have demonstrated that LLMs can overfit to sensitive information within training datasets, potentially resulting in the unintended exposure of private data during interactions through APIs, browser extensions, or other interfaces [17, 23, 24, 21, 49, 37]. Such risks are amplified by the susceptibility of LLMs to various attack vectors, including model inversion attacks (the reconstruction of sensitive data from model outputs), adversarial attacks such as prompt injection, and data poisoning attacks involving the introduction of malicious data into training sets [53, 6, 55]. To mitigate these risks, prior research has investigated privacy-preserving methods such as differential privacy, federated learning, and secure multi-party computation aimed at reducing the exposure of sensitive user information [52]. Despite these advancements, critical gaps remain to prevent LLM models from learning sensitive attributes during the interaction phase. Staab et al. highlight that even limited data from users’ online activities on platforms like Reddit or Twitter enables an LLM to predict private attributes with 85% top-1 accuracy [45].
Figure 1: Our Threat model
Comparison with prior works. Past research has primarily sought to study privacy-compromising data collection by extensions and analyze associated security vulnerabilities to protect sensitive user information from malicious actors. However, GenAI browser extensions integrated with modern LLMs introduce new risks. These include the potential exposure of user inputs, interactions, and browsing behavior to LLMs, which are highly efficient at inferring sensitive user attributes with high precision from seemingly benign interactions[45]. Our research builds on the findings of Staab et al. [45] by examining the compounded privacy risks of combining LLMs with browser extension capabilities, an area that has been largely overlooked. Our work is the first one focused at bridging this gap by performing a systematic audit of GenAI browser assistants. Additionally, past efforts have focused on analyzing leakage and sharing of data from user’s browser, without understanding how that data is used. GenAI browser assistants provide a unique vantage point to not just look at data leakage, but to also understand profiling and personalization capabilities based on the collected web-scale data. We investigate whether continuous tracking of user activities within their browser can help models associate memory with a user’s personal and behavioral data by periodically updating the context to personalize responses. Being a fairly new and developing space, it is important and timely to look at such unique risks posed by these assistants. We design a novel experimental framework to audit tracking, profiling, and personalization under 3 major functionalities – search, browse, and summarize – offered by 10 most popular search-based GenAI browser assistants.
4 Threat Model
Our threat model describes online users using GenAI browser assistants to personalize their browsing experience. The browser extension of the assistant installed onto the user’s browser is assumed to be the adversary while the user is considered the victim. The adversary operates in context of an extension providing it access to everything that the user does on the web. The primary goal of the adversary is to provide personalized responses to user’s queries. However, in order to achieve its primary goal, the adversary is assumed to engage in tracking victim’s browsing activities – by leveraging its presence as an extension. Thus user tracking is considered the secondary (or implicit) goal of the adversary. Adversary may share the collected information from the victim’s browser – either to its own servers or with third-party servers.
Figure 1 depicts our threat model. The victim can interact with the GenAI browser assistants in a variety of different ways – for example, making a Google search, summarizing the browsed pages, chatting with the assistant about the page, etc. Any such activity that involve a user’s interaction with the GenAI assistant is assumed to be a user’s query. Upon receiving a user query, assistant sends it along with metadata to its own server. The metadata may constitute user-, device-, webpage-, or browsing-specific information that serves as a “context” to generate a personalized response. GenAI assistant’s server may either generate a personalized response in-house or share the user’s query and context to an open-source LLM model using their API to generate a response. Finally, the personalized response is sent to the user’s browser and displayed to the victim
The GenAI browser assistant may also set and share first-party cookies from the victim’s browser to its own server. These cookies could further be shared with respective tracking platforms to either learn more about the user (i.e., analytics) or create custom audiences to re-target them with personalized ads on those third-party platforms. To fulfill its primary goal, adversary is assumed to continuously build and refine a user’s profile using the collected context. We consider LLM model’s server or third-party trackers to be beneficiaries as they may benefit from the data provided by the GenAI-browser assistant but are not actively involved in explicitly collecting information related to the victim.
5 Methodology
In this section we introduce our novel auditing framework for GenAI browser assistants to audit user tracking, profiling, and personalization as depicted in Figure 2. We make our crawling and analysis framework available at https://anonymous.4open.science/r/gen-ai-privacy-audit-1C72.
5.1 Selection of GenAI Browser Assistants
Table 1: Overview of studied AI browser assistants sorted by their popularity.
Legend: Personal Data : Personally Identifiable Information (PII), Personal Communications (PC), Financial Information (FI).
Web Data : User Activity (UA), Web History (WH), Website Content (WC), Location (LOC). No Data : No Data collected.
ExtensionName InstallCounts Supported Model(s) DefaultModel Invocation Mode Response Mode DataDisclosures SDK Version
Sider: ChatGPT Sidebar 4M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] sider Automatic Server-side PII WC 4.35.0
Monica - Your AI Copilot 2M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Mixed Server-side PII UA PC FI 7.6.0
ChatGPT for Google 2M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Mixed Client-side PII UA PC FI 5.5.1
Merlin Ask AI 1M [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o Mixed Server-side PII LOC 7.3.2
MaxAI: Chat with Webpage 800K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] [Uncaptioned image] gpt-4o-mini Manual Server-side PII UA 6.7.1
Perplexity - AI Companion 500K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] perplexity Manual Server-side No Data 1.0.21
HARPA AI 400K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] harpa-v1-smart Manual Server-side PII UA WH WC 9.6.2
Wiseone - AI Copilot 90K [Uncaptioned image][Uncaptioned image] gpt-4o Manual Server-side PII WC 1.7.2
TinaMind - AI Assistant 50K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] gemini-1.5-pro Manual Server-side PII UA PC 2.14.2
Copilot: AI Assistant 30K [Uncaptioned image][Uncaptioned image] [Uncaptioned image] gpt-4o-mini Automatic Server-side PII 1.5.73
Among the three categories of GenAI browser assistants described in Section 2, we focus on search-based assistants. This is due to several reasons: (1) Querying a search engine is relevant to querying an LLM model, allowing a natural integration of LLM-generated responses to a user’s online searches. (2) It has been argued that search engines and LLMs would co-exist due to their unique individual capabilities to incorporate recent information and generate context-aware responses, respectively [32]. (3) Although popularly referred to as ‘search’ based assistants, their capabilities are not restricted to search-only and they provide functionalities that allow their usage on any webpage as opposed to other categories of assistants that have limited scope and functionalities. (4) Combination of continued access to a user’s entire web browsing alongside generative capabilities raises privacy concerns, making them perfect candidates for our audit.
Our audit focuses on the most widely used web browser – Google Chrome. We survey various extensions from the Chrome Web Store to identify "AI-based search assistants" that offer the following three fundamental capabilities. (1) direct integration with browser search (i.e., Google search) via a sidebar response; (2) interactive sidebar functionality to chat with any webpage; and (3) ability to summarize webpage content. We select the top 10 most popular extensions based on number of installations that satisfy these requirements. The details about the selected assistants are shown in Table 1.
5.2 Crawling Infrastructure
To audit data collection and tracking practices of these browser assistants, it is important to examine network traffic activity emerging from the browser with the installed extension. We use a dedicated laptop to carry out our auditing experiments. The laptop is connected to a wireless Wi-Fi network to access the Internet and has Chrome browser installed. To intercept and log the decrypted network traffic while browsing the web, we install and configure Mitmproxy [39] – an open-source HTTPS proxy – on the device following official guidelines [19]. During the proxy configuration phase, Mitmproxy certificate is imported into Chrome’s Trusted Root Certification Authorities, allowing Mitmproxy to intercept and decrypt HTTPS traffic without triggering security warnings. The proxy server in this set-up acts as a man-in-the-middle between the laptop and the rest of the Internet as depicted in Figure 2, allowing us to capture all incoming as well as outgoing real-time web traffic while conducting our experiments – including requests, responses, payloads, and stored identifiers like cookies.
To conduct different experiments described in Sections 5.3 and 5.4 – each time, we first initialize a new Mitmproxy instance on localhost with mitmweb using command line interface (CLI). Next, we launch Chrome browser from CLI by specifying a new custom directory for user data to open a fresh browser instance with a new profile each time (Step ). This ensures a clean browsing session with no past history associated with the profile, resulting in independence across different experiments. Next, we visit the Chrome web store, search for the browser assistant, and install its extension into the browser instance (Step ). All assistants in our study, except Perplexity, require a user login to use it. Perplexity can be used with or without logging-into the account. To keep the profiles isolated across different experimental scenarios, we sign-up for a new user account each time using a distinct temporary email provided by various online services (Step ). However, some assistants doesn’t allow the use of a temporary email address – restricting it to either a Google or an Apple account. In such cases, we re-use an experimental Google account as Google prohibits creation of unlimited accounts tied to the same phone number. To avoid contamination across different experimental scenarios audited for a given extension, before the start of each experiment, we manually open a fresh browser instance to delete all Google activity and account-associated history. We also delete chat history and account-associated memory with the browser assistant’s account. We acknowledge that the browser assistant may store a server-side mapping of the user’s account-related activity data – which may or may not be deleted when the user performs client-side deletion. This is a limitation of our methodology. Next, we conduct the experiments described in the following sections. Finally, at the end of each experiment, web traffic from the entire browsing session is stored into a .flow file. We describe analysis of .flow files in Section 5.5.
Figure 2: Experimental design of our auditing framework for AI browser assistants.
5.3 Auditing User Tracking
When navigating through the web with an installed GenAI browser assistant, a user may visit a mix of public as well as private spaces online. In these spaces, a user may use the assistant without realizing the consequences of doing so. For instance, a user browsing through their health records may inadvertently end-up sharing sensitive information with the use of browser assistant. It is important to understand if the assistants are responsible and private with in-built safeguards to avoid exfiltration of user’s details or do they freely collect, share, and store information about the user at all times.
Experimentation. To understand implicit data collection and sharing, we identify 20 commonly-browsed website categories – 10 public spaces and 10 private spaces as listed in Table 6. For each category, we select one popular website to experiment with. All websites in private space are authenticated and contain personal or sensitive information that a user would not want to be collected, stored or shared. We experiment each website for every extension, totaling to a set of 200 experiments. Using our crawling infrastructure, we visit the website after logging-into the browser assistant’s account. Next, if it is a private space website, we login to the test website using (our own) personal account. If applicable, we navigate through the website by clicking on specific links that land us to the webpage containing private information about the user. For example, we navigate to ‘Medical Records’ tab and click on a specific visit record to open a detailed medical record associated with some visit. To avoid inconsistencies and ensure comparability, we always perform the same set of actions (if any) for a given website across different experiments. Once the website is loaded and navigations are complete, we invoke the sidebar interface and perform summarization of the webpage using its ‘summary’ feature (Step ). Finally, we ask a follow-up question (see Table 6 in appendix) to explicitly seek some information displayed on the webpage (Step ). This is to understand if the assistant recognizes the page content to be sensitive, problematic, copyrighted, or personal to avoid providing the requested information or not. Responses from the browser assistants as well as .flow files are stored for analysis described in Section 5.5.
5.4 Auditing Profiling and Personalization
As defined in Section 1, profiling involves collecting data pertaining to the users to draw inferences about them. However, using such a holistic view inferred from the profiled information to create individualized experience is personalization. It is important to understand if GenAI browser assistants associate the collected information about the user to their profile that it can recall to answer specific questions by the user, or to personalize its responses.
Prompting Framework. To evaluate profiling and personalization in GenAI browser assistants with respect to RQ3, we propose a novel prompting framework as depicted in Figure 2. The profile we develop represents – “A rich millennial male from southern California who is interested in equestrian activities”. We formulate 5 Train-Test prompt pairs – each explicitly leaking some attribute about the user in the train prompt and then testing for profiling based on the leaked information via a test prompt. The user’s attribute that we leak and test for in chronological order of prompts include location, age, gender, wealth, and interest. After the last test prompt, we re-ask all the 5 test prompts combined together as a single “profiling prompt” to the GenAI browser assistant. This is followed by “personalization prompt” where we simply ask the assistant to suggest Top 3 activities that the user would likely include in their itinerary of a vacation, based on what it has learnt about the user. The profiling and personalization prompts are first asked in the same tab where rest of the training or testing of individual attributes occur to understand its in-context behaviour. Next, we also open a new browser tab within the same session and re-ask them to understand its out-of-context behaviour. Except personalization and train prompts, we condition all responses to 5 test prompts and the profiling prompt to only output a binary response – "Yes" or "No" to avoid any subjectivity. This is achieved by adding a meta (system) prompt. Personalization prompt is reported to personalize responses (i.e., "Yes") if one of the suggested activities clearly lists atleast one ‘equestrian activity’ (i.e., the user’s leaked interest). We explicitly choose a niche interest for our profile to avoid matches with generic activities that an LLM model might suggest such as ‘hiking’ for instance. Section 8.2 in appendix lists all prompts used in this study.
Experimentation. Amongst different features provided by search-based GenAI browser assistants, we specifically test three most useful features that were observed to be common across all 10 extensions – search-based integration, webpage chat, and webpage summary. These are studied under the following four scenarios for each extension (totaling to 40 experiments) as described in Figure 2:
Control. In control, we simply ask each test prompt (without train prompts) followed by profiling and personalization prompts – each in a new tab.
Search. In search scenario, train prompts are entered into Google search as a normal search query. When Google search is performed, it may result in an automatic response generation from the GenAI assistant as well, which is displayed to the user as a sidebar response. In cases where automatic responses were not generated, the train prompt is explicitly asked to the GenAI assistant in the sidebar chat. Next, the test prompt corresponding to the train prompt is asked as a follow-up question in the sidebar chat. The same steps are repeated for each train-test prompt pair within the same browser tab. At the end, profiling and personalization prompts are asked both – in-context and out-of-context.
Browse. In this scenario, no interaction occurs during the training phase. Training phase involves browsing through 10 webpages – 2 pages per leaked attribute. These webpages (listed in Table 5 in appendix) are selected such that they leak different attributes about the user. Other than scrolling a page to explore its content, we also click on internal links to emulate normal browsing. The set of clicks are fixed and remains the same across different experiments involving the same website. Our hypothesis is that if GenAI browser assistant collects information about the user’s visit to different webpages, then it may use it to infer user’s attributes. The experiment ends by asking 5 test prompts along with profiling and personalization prompts sequentially within the same tab in the sidebar chat. Profiling and personalization prompt are also repeated out-of-context.
Summarize. We use “summarize webpage” feature of GenAI browser assistants to understand if usage of this feature aids the browser assistant in building a profile about the user and later on use it to personalize its responses. We visit the same 10 webpages as before, performing the same interactions as in ‘Browse’. Additionally, we also summarize each page. In this scenario, summarization acts as training. Finally, the same prompting as in ‘browse’ is performed.
We repeat each prompting experiment in our study 3 times and report the majority response to account for variability in the output due to probabilistic nature of GenAI systems.
5.5 Semi-automated Analysis
In this section, we discuss our approach to perform network traffic analysis based on the web traffic data collected in the form of .flow files during the experiments. We analyse .flow files corresponding to the most representative (i.e., majority) response. We use Python module mitmproxy.io to programmatically parse the network flows and extract details about the contacted endpoints, request and response headers, payloads, and responses. We identify each request to be first-party if its domain matches the domain of the browser assistant’s website, and third-party otherwise. A flow is also classified as either a foreground flow or a background flow. Background flows represent extension traffic emerging from the background service worker of the browser assistant. They are identified based on the value of a request header – origin – value of the form chrome-extension://<extensiod-id> represents traffic from an extension, where extension-id can be obtained from extension’s page on Chrome web store. The automated code extracts the flows from different experiments related to a given extension along with the above details in the form of a CSV file. We perform detailed analysis of these files to extract user, device, webpage or browsing related information from the payload.
6 Results
6.1 Architecture of GenAI browser assistants
This section discusses our findings related to RQ1, providing a deeper understanding on how GenAI browser assistants are designed. We qualitatively analyze network traffic activity to learn the following architectural differences across 10 assistants: (1) backend model, (2) response architecture, (3) context restrictions, and (4) response variability.
Backend model. Since GenAI browser assistants function as wrappers on top of the open source LLM models, we observe that they provide support for most of the popular open-source models – OpenAI’s chatGPT [36], Google’s Gemini [16], Anthropic’s Claude AI [4], and Meta’s Llama [26] as shown in Table 1. For each extension, the request to fetch the response also contained the model information. We notice gpt-4o-mini to be the most popularly used default model due to its cost-efficiency trade-off for assistant developers as well as for users over other assistants. Some extensions used different models for different user activities. For example, ChatGPT for Google (CFG) used gpt-4o-mini for Google search triggered automatic response generation while it used gpt-3.5-turbo when a user explicitly opened a sidebar chat to ask some question. Sider, Perplexity, and Harpa use custom LLM models, making it difficult to understand how its design differs from other open-source models.
Table 2: Data collection and exfiltration behavior of assistants in public and private online spaces of a user. Exfiltration legend:
Full Webpage : Page text, title, location, hyperlinks. Server-fetch Webpage : Page title, location, server-fetched file’s upload location. Plain Webpage : Page text, title, location. Partial Webpage : Partial content or missing details. Response legend: ✔: Response with Relevant Details. ✔: Missing some details in Response. : Response restricted. ✘: No response generated.
Category WebPage Sider Monica CFG Merlin MaxAI Perplexity Harpa Wiseone TinaMind Copilot Public Spaces News Platforms cnn.com Open Forums reddit.com Informative Articles wikipedia.org E-commerce Website amazon.com Sports Websites espn.com Travel Platforms expedia.com User-generated Media youtube.com Kids Website nickjr.com Misinformation Website infowars.com Violence Material guns.com Private Spaces Health Portal university health portal Email Account mail.google.com Social Media Platform facebook.com Adult Content pornhub.com Online Streaming Service netflix.com Government Website irs.gov Dating Service tinder.com Financial Service chase.com Educational Platform canvas.instructure.com Messaging Platform slack.com
Response architecture. Browser assistants may contact open-source GenAI models to generate responses either from client-side or server-side. Figure 1 depicts the most common case we observed in our study, where 9 out of 10 extensions operate server-side. This means that first, the user’s selected model (or default in case of no selection) is shared along with the user’s query and other metadata to the browser assistant’s server. Based on the choice of model, assistant’s server would then call API of the respective LLM model to direct the user’s request to. The personalized response is finally displayed to the user. In case of CFG, response generation is initiated from the client-side (see Table 1). To use CFG browser assistant, the user needs to link their OpenAI’s chatgpt.com account with CFG. As a result, all queries from interaction with CFG are directly shared with chatgpt.com to fetch the response. We also analyze if browser assistants always monitor user’s activities (passively) or if they perform monitoring only when user explicitly interacts with the assistant (actively). Here we refer to activities that do not exist as features in the browser assistant, such as scrolling, clicks, navigation. etc. We found it to be always “active” for all 10 extensions, suggesting that they do not passively monitor user’s non-extension related activities. In regards to extension-related activities like search, we found that some assistants such as Sider and Copilot automatically invoke response generation when a user submits a search query in the search engine. 5/10 extensions require manual invocation of the assistant to obtain a response while remaining 3 work in a mixed fashion.
Context restrictions. Next, we compare the interdependence of context used by the browser assistant under different scenarios. We look at two sets of context – sidebar response versus popup sidechat and context across page navigations. In the former case, during our experiments, we analyse dependency of context from the response generated by the assistant in the empty sidebar space of Google search results page with the context of the sidebar chat that pops up upon clicking the extension icon. We did a simple experiment, where we asked in the Google search ‘I like apples’ (Q1) and ensured that the sidebar response shows up. Next, we opened the sidebar popup chat and asked ‘Do I like apples?’ (Q2) – if it responds ‘Yes’ then that suggests the two contexts are interpreted as dependent, otherwise independent. We observed that 4 assistants – ChatGPT for Google, Max AI, TinaMind, and Copilot maintained the same context while the remaining 6 maintained separate contexts, suggesting stronger privacy. In the second case, we ask Q1 in the sidebar popup chat when visiting some webpage; next we navigate to a different webpage and ask Q2 to understand if it preserved the context or not. We observe that context was not reset only for 2/10 assistants – ChatGPT for Google and WiseOne, suggesting their capability to remember contexts across multiple websites as user browses through the web. We found Perplexity to be the most private assistant as it explicitly displayed the message stating ‘I do not have the ability to recall previous interactions or questions. Each session or question is treated independently for privacy reasons.’. We further analyze what data forms the necessary context for different assistants in more detail in Section 6.2
Response Variability. It is important to understand that despite different browser assistants using the same model to perform the user’s query, the outputs returned to the users could be completely different. Several factors play a role in this. We discuss two of the most important ones below – Temperature and System prompts.
Temperature. First, as aforementioned, output of generative models is probabilistic and is influenced by the temperature parameter. For example, OpenAI’s ChatGPT provides capability to set temperature in the range 0 to 1 where lower values provide more deterministic and focused outputs while higher values generate more creative or random responses [35]. Different browser assistants may set and use different values of temperature, resulting in differences in outputs even when they use the same model. Moreover, even the same browser assistant may dynamically set temperature for different user queries. For example, when asked to a browser assistant: “What do you know about me?” – it may query OpenAI’s API with a higher temperature value to not generate an extremely precise (or invasive) output and rather keep it open-ended. However, if the same user asks the same browser assistant: “What is the distance between New York and Washington D.C.?”, the assistant may use a smaller value of temperature to provide a deterministic output.
System Prompts. Second, system prompts are provided to the LLM models along with the user query for various reasons such as to enhance task-specific accuracy, provide contextual guidance, prevent unwanted output, ensure output formatting, and mitigate vulnerabilities such as jailbreaking [48]. During the network traffic analysis, we discovered various system prompts being empoloyed by different browser assistants as listed in Section 8.3 in appendix. We observe some assistants to use a tighter system prompt to answer a user’s query more accurately than others. For example, Sider’s system prompt to answer a question simply states to “use simple and clear language” whereas Harpa’s system prompt explicity states to “NEVER fabricate, infer, or guess information. Do not hallucinate links. Be to the point”. In the same prompt, it can be seen that Harpa includes $\{\{user\_info\}\}$ as a context along with user’s query to generate a response, suggesting that it shares user details with open-source LLM models freely with each query. Moreover, Harpa’s system prompt states “Please ignore all previous instructions” to prevent any jail-breaking attempts through cleverly-crafted user queries. TinaMind’s system prompts are observed to associate language preference, day, date, timestamp, and timezone information with each query in order to interpret the user’s location-dependent queries as per the applicable region to produce a more relevant response.
6.2 User tracking by GenAI browser assistants
To systematically audit user tracking by different browser assistants, in Section 6.2.1, we first perform analysis of implicit collection and sharing of user’s data as user visits public and private spaces online. We refer to it as ‘implicit’ as the user does not want to actually share their personal information with the GenAI assistant. However, the user may inadvertently let the assistant access it. Here, we analyse tracking of user data in context of the webpage content that the user is visiting. Next, in Section 6.2.2, we evaluate explicit data collection and sharing when using different functionalities offered by assistants, where we use the prompting framework to explicitly leak user attributes and understand tracking of all kinds of data associated with the user.
6.2.1 Implicit collection and sharing of user data
In this section, we discuss the results related to implicit user tracking to see if and when do browser assistants collect data. We followed the methodology described in Section 5.3 to visit different websites in public and private spaces, summarize the page content, and ask a follow-up question. Table 2 showcases results from running a total of 200 experiments across 20 websites and 10 assistants.
The primary goal of the analysis is to understand what kind of webpages are vulnerable to collection of page content related data by browser assistants. We particularly focus on page content due to its direct impact on user’s privacy. This is because if the visited webpages happen to be one of the private spaces to the user, then it puts their private data at a risk of getting collected. Besides user privacy, data collection of problematic content can result in browser assistant providing harmful responses to the user, while collection of copyrighted content may have regulatory implications.
As part of the page data, we consider page title, page text, page location and any embedded links on the page. The cells in red suggests most egregious data collection, where complete webpage details were collected, while cells in green correspond to inefficient collection of pages (e.g., with missing details). We observe that Harpa collects full DOM in all 20 online spaces, while TinaMind is unable to extract full webpage details across all the experiments. This is likely due to two reasons – one, TinaMind only focuses at capturing data that is in the viewport of the user, failing to capture rest of the page and two, its page extraction mechanism is less aggressive. The latter inference is based on the fact that it performed incomplete extraction even from the user’s viewport content, often failing to reproduce the numbers present on the webpage in its responses.
On their Chrome webstore page, Harpa clearly states “we do not collect or sell userdata” – contradictorily, we observe it to collect 100% of health records, student academic data, as well as user’s personal messages on messaging platforms. As mentioned in Section 6.1, Harpa’s system prompt also included the user’s name and location in plain text, for example: “My name is John Gabb. I am in London, United Kingdom of Great Britain and Northern Ireland. Please answer in English…” This demonstrates clear violation of their own privacy practices.
Browser assistants can easily employ practices that can allow them to determine if a page contains sensitive or private data or not. For instance, when summarize feature is used with Perplexity, it shares the page URL with its own server and performs a server-side fetch of the webpage. This is inferred based on the file upload location received in response to the shared page location. Therefore, if the webpage belongs to a user’s private authenticated space, Perplexity’s server will never have access to the user’s personal data. However, it will still be able to provide responses for public spaces. In contrast, Wiseone employs a different approach by obtaining explicit user consent for each new website where the user wishes to utilize its functionalities.
Merlin, MaxAI, ChatGPT for Google and Monica were all able to extract webpage contents for all 20 scenarios. Merlin was discovered to be the only assistant that recorded even the contents of the forms on webpages as opposed to all other browser assistants – that did not collect form data. For example, it was able to collect the user’s “Social Security Number (SSN)” entered in a form on IRS refund portal. Email addresses as well as the full email thread were also collected by these assistants. One of the most shocking findings was that GenAI browser assistants were freely able to collect and share data to their own servers on authenticated health portals. They were able to answer follow-up questions ranging from patient details to entire medical history. Collection of PHI without appropriate user consent is in clear violation of HIPAA [14]. Moreover, student’s academic records including assessment scores, exam performances, overall grades – were all collected and shared with browser assistant’s servers demonstrating violation of FERPA [34] that aims to protect these attributes for a student.
Overall, it can be observed that responses with missing details are more prominent in the lower half of the table suggesting that private spaces are handled with care by most extensions. However, this is insufficient since they are still collecting partial data that is private to the users. Adult content, online streaming services, dating platforms, and financial services – were unanimously responded inefficiently by all the assistants. Titles displayed on Netflix homepage, preferences set on Tinder and transaction amounts as well as last four digits of card number on Chase were either inaccurately output or were missing from the output of different assistants. This is because we observed that the assistants missed capturing shadow elements from DOM of the webpage. On the brighter side, Perplexity and Wiseone actively suppressed response generation on adult pages due to explicit content stating “I apologize, but I cannot provide information about or analyze pornographic content. This type of material is not appropriate for me to discuss. Perhaps I could assist you with a different, non-explicit topic instead?”. However, both of them still recommended links to pornographic content in follow-up suggestions, showing inefficiencies in their model.
Table 3: Data sharing across extensions. Legend: = First-party sharing, = Third-party sharing, = Both, = None.
Scenario Page Location Page Content GSearch Results User’s Query Chat ID Chat History User Details Device Details Query Timestamp Timezone Referrer Cookies User Agent Local Storage
Sider Search
Browse
Summarize
Monica Search
Browse
Summarize
CFG Search
Browse
Summarize
Merlin Search
Browse
Summarize
MaxAI Search
Browse
Summarize
Perplexity Search
Browse
Summarize
Harpa Search
Browse
Summarize
Wiseone Search
Browse
Summarize
TinaMind Search
Browse
Summarize
Copilot Search
Browse
Summarize
6.2.2 Explicit collection and sharing of user data
Having looked at what online spaces are vulnerable to data collection, we now focus on understanding what attributes are explicitly collected about the user, device, or page when different features provided by the browser assistant are used by the user – namely, search, browse and summarize as shown in Table 3. We look at the requests, payloads, and headers to identify different attributes that are collected and shared with either first-party server of the browser assistant, third-party servers, both or none during different scenarios.
Page Data. Distinct from the implicit data collection behaviour discussed in Section 6.2.1, we observe that besides their first-party servers, Harpa and MaxAI also share page location with the third-parties. Specifically, they both share it with a third-party pixel tracking company – api.mixpanel.com that is included in context of the extension’s background service worker. Mixpanel offers analytical as well as session replay [30] capabilities. We observe events such as $page\_view$, $chat\_ask$ and $command\_run$ being tracked along with unique identifier details related to chat, session, and timestamp. Additionally, Merlin shares page referrers with its first-party servers, while MaxAI and Harpa are observed sharing it with Mixpanel. An important thing to note is that MaxAI-injected content script also included Mixpanel JS in foreground, resulting in event tracking by the endpoint api-js.mixpanel.com. During search experiments, we see that Monica, Sider, and Merlin are the only extensions that collect and share Google search results displayed to the user along with the user’s query with their respective first-party servers. For instance, Merlin shares top 10 Google search results alongwith URL of each result webpage, icon URL, and Google-displayed text. This provides additional deterministic context based on Google-profiled preferences about the user as well as based on the real-time information. This aids browser assistants to fine-tune their responses more accurately, which purely based on probabilistic LLM-generated response could have become less relevant as discussed in Section 2.1 and Section 6.1.
User’s Data. Now, we look at collection and sharing of user’s search query, chat details, and user-related identifiers with different endpoints. Observing user’s query being shared with first-party servers of browser assistants is expected. However, surprisingly, we found Merlin’s background service worker to include Google Analytics tracking script. As a result, user’s raw query was also shared with google-analytics.com endpoint. In case of CFG, we observed the query being shared with chatgpt.com to fetch a response to the user’s prompt. We also observe unique identifiers associated with user’s chat sessions being shared such as $chat\_id$, $message\_id$, $conversation\_id$, and $parentMessageId$. Such chat identifiers were shared with all first-party servers, except Perplexity and Wiseone. In case of TinaMind, chat identifiers were shared with analytics.google.com along with the user identifiers. However, Sider and Merlin were observed sharing similar user identifiers with google-analytics.com. An important distinction is that sharing data with google-analytics.com allows tracking the user for the purpose of analytics. However, sharing data with analytics.google.com allows joining user’s identity across Google’s domain google.com with shared cookies. Browser assistant developers can create custom audiences based on the query terms or chat identifiers to (re-)target users with ads across Google properties such as mail.google.com for instance. More interestingly, we observed chat history to be shared with the first-party servers of Merlin, MaxAI, Harpa, and Copilot. Harpa and Copilot maintained the entire chat history of the user since their first conversation in the background service worker’s IndexedDB storage and the full history was shared with every new query to provide complete context. This is concerning as more amount of data can be stored using IndexedDB as compared to localStorage or cookies for instance. Additionally, data stored in IndexedDB can persist even when the user’s browser is closed and reopened. Perplexity stored states using what it referred to as rwToken in localStorage. In terms of cookies, we observe many extensions setting first-party cookies like _fbp (Facebook), _ga (Google analytics), _clk (Clarity), etc. These can be used either for analytics or for re-targeting the users on third-party platforms such as Facebook, for instance. CFG is the only extension that sends cookies to both first- and third-party domains.
Thus, third-party data collection and sharing is more concerning than first-party as it allows linking and targeting of user across multiple websites. Figure 3 summarizes all third-party sharing observed by us. Moreover, first-party data collection and sharing suggests potential of profiling and personalization as we discuss in the next section.
6.3 Profiling and Personalization
In this section, we aim to understand if the data collected and shared by the first-party servers of browser assistants is stored and used for profiling of the user. Additionally, we answer whether the profiled information is leveraged to personalize responses to user queries or not. We followed our novel prompting framework described in Section 5.4 to invoke profiling and personalization. Table 4 shows our results. We observe ChatGPT For Google, Copilot, Monica, and Sider
Table 4: Extensions’ Personalization and Profiling Results. * corresponds to logged-out state.
Legend:  signifies profiling or personalization shown in response.  ×  signifies no profiling or personalization shown in response.
Test Prompts Location Age Gender Wealth Interests Profiling Personalization Profiling Personalization (Test 1) (Test 2) (Test 3) (Test 4) (Test 5) (In-context) (In-context) (Out-of-Context) (Out-of-Context) Expected Results Sider Control × × × × × × × × × Search × Browse × × × × × × × × × Summarize × × × × × × × × × Monica Control × × × × × × × × × Search Browse × × × × × × × × × Summarize × × × × × × × × × CFG Control × × × × × × × × × × × × × × × × Search Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × Merlin Control × × × × × × × × × × × × × × × × Search × × × × × × × Browse × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × Max AI Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × Browse × × × × × × × × × Summarize × × × × × × × × × × × × × × × × Perplexity Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × Control* × × × × × × × × × × × × × × × × × Search* × × × × × × × × × × × × × × × × × Browse* × × × × × × × × × × × × × × × × × Summarize* × × × × × × × × × × × × × × × × Harpa Control × × × × × × × × × × × × × × × × × Search × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × Wiseone Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × TinaMind Control × × × × × × × × × × × × × × × × × Search × × × × × × × × × × × × × × × × × Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × × × × × × × × × × CoPilot Control × × × × × × × × × × × × × × × × × Search Browse × × × × × × × × × × × × × × × × × Summarize × × × × × × × × ×
to demonstrate expected behaviour for the search scenario where they were able to profile the user based on each leaked attribute regarding location, age, gender, wealth and interests. Additionally, it also personalized user responses both in-context and out-of-context. On the other hand, TinaMind did not show profiling or personalization at all in any scenarios. On a similar note, Perplexity also did not show any profiling, responding to personalization prompt with “I apologize, but I don’t have any prior information about you or your preferences to base personalized recommendations on. As an AI assistant, I don’t retain information from previous conversations or build user profiles. Each interaction starts fresh.”. Moreover, Monica and Sider were observed to not fully profile all leaked attributes about the user in non-search scenarios – only age and interest attributes show consistent profiling in action across scenarios. It is interesting to see that the profiled information transcend beyond context restrictions discussed in Sections 6.1 for Monica and Sider, suggesting that the browser assistant is likely maintaining a profile for the user on server-side to be able to personalize its responses in-context as well as out-of-context in all cases. Surprisingly, no extensions except Monica and Sider, show any profiling or personalization for the contorl scenario. The fact that these two show a positive response in control, suggests that these browser assistants might be associating even the asked questions to be reflective of the corresponding leaked profile.
Harpa demonstrates profiling for all 5 attributes for search and in-context personalization. However, it fails to retain the profile out-of-context. Merlin and MaxAI demonstrate similar profiling behaviour for interest attribute across both search and browse scenarios, resulting in in-context personalization that doesn’t similarly translate to out-of-context. Behaviour of Merlin and MaxAI shows unpredictability – for instance, Merlin search shows profiling at each step of the test prompting. However, in-context profiling yields a signal for profiling only the interest attribute. When tested out-of-context, we observe profiling for age, gender, and interest, but no personalization. We believe the same reasoning could apply here as discussed previously – dependency on the value temperature parameter used by Merlin. Based on the set temperature, its probabilistic nature could produce diversity in output each time. Many extensions result in the output “I’m sorry, but I cannot answer those questions with "YES" or "NO" as I do not have any information about you.”. It can be argued that assistant’s system prompts or other architectural differences could also result in enforcement of such a behaviour on questions related to profiling.
Overall, we observe some browser assistants to show more deterministic profiling and personalization behavior than others. We observe search-based profiling to be the strongest for most extensions, demonstrating how user’s data tracked through their searches can easily lead to their profiling for personalization of responses by GenAI browser assistants.
7 Conclusion
In this paper, we systematically evaluated the 10 most popular GenAI-based browser assistants to investigate their architectural design, privacy risks, and profiling and personalization capabilities. Our findings indicate that: (1) 90% of the browser assistants rely on server-side processing, contacting third-party LLM models from their first-party servers by transmitting user queries and metadata for response generation. Notably, two of the extensions (Sider and Copilot) are auto-triggered even when users perform a query on the browser search engine. 60% of the GenAI browser assistants isolate the contexts of each query. However, ChatGPT for Google and Wiseone maintain context across page navigations. (2) Two browser assistants (Harpa and Copilot) collect the full DOMs of user-visited pages, while others gather varying levels of private data from webpage content, such as medical records, student’s academic records, or query-based Google’s search results. Additionally, Harpa and MaxAI share page locations and referrers with third-party tracking services. Merlin was discovered to be the only assistant that recorded the contents of web forms and managed to collect and share SSN as part of a webform. Moreover, some assistants, such as Sider, Merlin, and TinaMind, share chat identifiers and raw queries with Google Analytics, enabling potential tracking and retargeting on external platforms. (3) GenAI browser assistants demonstrate varying degrees of profiling and personalization based on user attributes such as location, age, gender, wealth, and interests. Browser assistants such as ChatGPT for Google, Copilot, Monica, and Sider perform extensive profiling across all five attributes, enabling both in-context and out-of-context personalization when search functionality is used. However, others, such as Perplexity and TinaMind, do not show strong signals of profiling or personalization. Overall, we observed Perplexity to be the most privacy-friendly while extensions such as Harpa, MaxAI, and Merlin were amongst the least.
As highlighted in our work, capabilities of GenAI assistants are tremendous in terms of the granular level of access that these models have to user’s personal or sensitive data. This capability underscores the urgent need for robust regulatory frameworks. Collection and sharing of medical records and student’s academic records by different browser assistants demonstrated non-compliance with the regulations such as HIPAA [14] and FERPA [34]. Browser assistants can easily incorporate a check to decide whether or not the webpage contains PII or sensitive information about the user via domain-based classification. Moreover, they could also adopt Perplexity’s approach to perform a server-side fetch of the webpage based on the URL to avoid collecting any personal data about the user or obtain explicit consent from the user on each webpage similar to Wiseone. In absence of any regulatory enforcement, most of these assistants do not have any mechanisms in-place to evaluate their privacy risks. We recommend policymakers to adopt a bottom-up approach to regulate GenAI assistants as they increasingly influence the future of web browsing and search engines. Privacy must be embedded into these systems by design, and developers must actively mitigate risks of data leakage, profiling, and unauthorized sharing.
As advancements in GenAI shape the future of web browsing, some envision an in-browser model powering an intelligent yet private web browser that can automate different tasks to simplify day-to-day browsing needs of a user. While others anticipate a paradigm shift [44] where traditional browsers and search engines may become obsolete. These transformations are currently being driven by advances in artificial general intelligence (AGI) that aims to automate workflows by integrating different specialized agents that work together to complete user’s task such as booking a dinner reservation in an end-to-end manner. Google is building towards the former vision and has already begun origin trials for Built-in AI [11] to allow developers of websites and webapps to leverage Gemini Nano through several APIs. On the other hand, Microsoft’s Recall [1] recently brought GenAI to user’s computer allowing them to record their screen and integrate recordings with GenAI locally. More recent developments such as Window.ai [2] and Screenpipe [18], have also been in similar directions as Recall, inspiring user-centric designs for managing GenAI interactions directly within the user’s device, reducing reliance on server-side processing and ensuring user privacy. As GenAI browser assistants increasingly shape this landscape, it is important that privacy considerations form the foundation of their design and deployment. Future works should delve deeper into improving transparency mechanisms for GenAI-based browser assistants. Collaboration between developers, researchers, and policymakers will be critical to ensure these tools empower users while safeguarding their rights and privacy. We believe our work will lay a foundational ground to inspire future privacy work in this space.
Ethics considerations
Our research involves visiting private online spaces such as personal health portal, social media platform, personal email, etc. to understand the risks posed by using GenAI browser assistants in such spaces. We did not recruit any participants for our study and used our own personal accounts to perform the experiments. We reached out to the Institutional Review Board (IRB), but they granted an exemption to our research since it involved the use of personal accounts and therefore did not fall under the definition of human subjects.
Our research involved creating and using accounts with various browser assistant providers to evaluate their services during web browsing experiments. To ensure ethical practices, we utilized temporary email accounts, created no more than one account per assistant per day, upper bounded to 12 accounts, and refrained from excessive account creation. Most of the experiments were conducted using the free token credits provided at sign-up, but to ensure consistency in the features tested across different assistants, we also purchased a monthly paid subscription for some extensions.
Although this study did not involve human subjects, we adhered to ethical principles aligned with those outlined in The Menlo Report. Specifically, we prioritized beneficence by minimizing potential harm, respected the autonomy and integrity of organizations, ensured fairness in the evaluation of different assistants, and remained compliant with applicable laws and public interests.